Leona owns her own business and works full time in the store without paying herself a salary. She has $20,000 of her own money invested in the store that she withdrew from her savings account, which earned 10 percent interest. She was offered a job last year making $28,000 per year, but turned it down. If Leona's accounting statements show revenues of $100,000 and accounting costs of $60,000, then Leona's

A. accounting profit is $20,000, and her economic profit is zero.

B. accounting profit is $40,000, and she is making an economic loss of $8,000.

C. accounting profit is $40,000, and her economic profit is $10,000.

D. accounting and economic profit is $40,000.
Answer #1

Answer

Accounting profit =revenue -accounting costs

=100000-60000

=$40000

Economic profit =accounting profit -implicit costs

implicit costs =salary of given up job + given up interest on the savings

=28000+20000*0.1

=30000

Economic profit =40000-30000=$10000

Option C